# Geographic Variation in

### v1

@@ 3.2 Mobility maps
We can compare our estimates to Heidrich (2017), who provides detailed national and regional estimates of intergenerational income mobility for the Swedish population born between 1968 and 1976. We consider more recent cohorts and our empirical specification differs in various aspects. In particular, we predict prime-age earnings based on all available earnings data of each individual, compared to using a conventional shorter-run average of observed earnings or income. Another important difference is that we consider variation on the finer municipality level, while Heidrich focuses on the broader definition of local labor market regions. [^19] One could debate which level of aggregation makes more sense. Child care, compulsory school and high schools are organized at the municipality level, and financed by municipal income taxes. Many other local amenities and services are also best distinguished at the municipal level. Local labor market regions are larger and defined based on commuting patterns. Different divisions are thus likely to capture differences in mechanisms underlying intergenerational transmission to different extents. The exact division is not key in our setting, as we primarily aim at making a set of conceptual and methodological points. Consistent with this and other differences in specification, we find slightly higher earnings persistence. Nevertheless, the spatial pattern appears largely similar. For example, Heidrich (2017) notes that several local labor markets in the south of Sweden and around Stockholm are doing particularly well in lifting children from lower income families, while absolute upward mobility is particularly low in municipalities in the Swedish inland located close to the Norwegian border. These patterns are also visible in our Appendix Figure A.5b. Given that we study different birth cohorts, this similarity also implies that mobility differences between regions are systematic and at least somewhat persistent, rather than just temporary.
@@ 3.3 Linear approximations
Regional comparisons of intergenerational mobility, such as those reported in Figures 1 and 2, are subject to two methodological concerns. The first is whether intergenerational relations can be sufficiently well summarized by simple summary measures, such as linear regression or correlation coefficients. Cholli and Durlauf (2022) note that theoretical models usually imply nonlinear relationships between parent and child status, which are often ignored in practice.
For illustration, Figure 3a plots the mean child earnings rank against the parental earnings rank at the national level. The rank-rank relation is linear over most of the distribution, but much steeper at the very bottom and top. [^20] Nybom and Stuhler (2017) document a similarly steep relation at the very bottom of the parental income distribution. The conditional expectation function (CEF) is therefore not fully captured by a linear regression, in contrast to the pattern for the U.S. as described by Chetty et al. (2014). However, it serves as a good approximation over most of the distribution, such as between the 10th and 90th percentiles. The expectation of child years of schooling conditional on father’s years of schooling is approximately linear for most of the distribution as well (see Appendix Figure A.6).
@@
While the CEF appears sufficiently linear on the national level, it does not necessarily follow that the conditional expectations are equally linear on the regional level. For illustration, Figure 3b plots the mean child earnings rank conditional on parent rank deciles for Sweden’s five largest municipalities. Upward mobility is substantially lower, and the rank-rank slope steeper, in Malmö (diamonds) compared to for example Uppsala (grey circles). However, apart from such differences for the bottom decile, the relationships are approximately linear in all cities; while it is still somewhat steeper at the bottom and the top of the parental earnings distribution, this pattern is less pronounced at the regional level. We provide the corresponding graphs for the grandfather-child relationship in Figures A.7a and A.7b in the Appendix. The region-specific multigenerational CEFs are approximately linear, and less steep at the bottom of the income distribution than their national-level counterpart. [^21] Further, we conducted a simple analysis for all municipalities by comparing the regression $R^{2}$ from municipality-specific linear and quadratic regressions. The quadratic model generates an $R^{2}$ that on average is only 2.7% larger than the mean $R^{2}$ from the linear model. This number drops to only 1.4% if we weight the averages by number of individuals rather than by municipality. Linear measures therefore appear well suited for cross-regional comparisons.
@@ 3.4 Sampling error
A second concern in fine-grained regional comparisons is sample size: Some of the variation across municipalities might not be “real” variation but reflect sampling error, in particular for smaller municipalities for which comparatively few parent-child pairs are observed. [^22] As we observe (almost) the entire population, our estimates do not suffer from traditional sampling error. However, the observed mobility for a given set of parent-child pairs can be interpreted as a random draw from the “potential” distribution for that municipality. While the mean number of father-child pairs per municipality for children born between 1981 and 1989 is 3,183 for education and 3,099 for earnings, in 57 of the municipalities (59 for earnings) we observe fewer than 1,000 pairs. [^23] Moreover, inference based on earnings ranks is complicated by the fact that each unit’s rank depends on which other units have been sampled (Mogstad et al., 2023). We also observe fewer grandfather-child pairs than father-child pairs, and especially so for earnings. Education is observed for all who have resided in Sweden at some point as adults, while earnings is only observed for those who at some point from 1968 have recorded taxable labor earnings above a certain lower amount (see Section 2). To probe the influence of sampling error, we implement two tests.
@@
Notes: Density plot of the estimated intergenerational correlations in years of schooling for father-child pairs across 160 “small” municipalities ($\leq$ 2,000 observations), 130 “large” municipalities ($>$ 2,000 observations), and corresponding “placebo” municipalities in which we randomly allocated pairs across municipalities.
@@
First, we split the sample into smaller ($\leq$ 2,000 observations) and larger municipalities ($>$ 2,000 observations), with the expectation that the influence of sampling error should be more limited for the latter. Figure 4 plots the resulting densities of the father-child correlation in years of schooling. The correlations vary nearly as much for large as for small municipalities (the standard deviations are 0.038 vs. 0.048). Figure A.8 in the Appendix shows corresponding distributions for earnings. [^24] Figure 4 shows that the educational correlations are on average substantially lower for small municipalities, a pattern that is possibly related to the greater homogeneity in educational attainment in smaller municipalities. And while the estimated coefficient in an OLS regression of the municipal parent-child correlation on sample size is large and significant, it loses significance if we control for the standard deviation of years of schooling. The difference in means between small and large municipalities for earnings ranks is smaller, but qualitatively similar (see Figure A.8 in the Appendix). Second, we compute “placebo correlations” by reshuffling parent-child pairs randomly across all 290 municipalities, such that variation in the placebo correlations solely reflects chance. Figure 4 shows that for both small and large municipalities, the resulting density is much narrower than the actual variation of intergenerational correlations across municipalities. Moreover, there is considerably less dispersion in the placebo distribution for large than for small municipalities – the standard deviation of the former is less than half of the one for the latter. [^25] Considering instead changes, the decrease in dispersion is larger for large than for small municipalities when comparing the placebo with the true distributions, with a drop in the std. dev. of about 70% compared to 45% for small municipalities. These differences between small and large municipalities are even more pronounced for earnings ranks. In particular, the std. dev. of the placebo distribution of earnings rank slopes of large municipalities is about 44% of the one for small ones.
@@
Still, some of the reported variation between municipalities reflects noise, in particular for small municipalities, and sampling error might explain much of the regional variation in studies that use less data than we do here. Sampling error may have a greater effect for earnings-based measures, as earnings are themselves subject to transitory noise (Mazumder, 2005). This is the case in our setting as well, even though we use comparatively high-quality earnings measures (see Appendix Figure A.8). Few studies have addressed this issue systematically. Rather than estimating separate regressions for each region, Heidrich (2017) estimates a multilevel model that explicitly accounts for the influence of sampling error on the variation of regional mobility estimates. This approach effectively “shrinks” the region-specific estimates to their mean, in particular for small regions with imprecise estimates. Risa (2019) implements a similar shrinkage procedure for mobility estimates for Norway. In addition, Heidrich (2017) reports only those estimates that are significantly differently from the mean, replacing all other region-specific estimates with that mean.
@@ 4 Are Regional Rankings Stable?
Our main focus is on whether intergenerational (parent-child) associations – which can be estimated in many settings – reflect similar underlying mobility processes as more data-demanding multigenerational associations. This analysis is complementary to a recent analysis by Deutscher and Mazumder (2023), who compare the ranking of Australian regions across different measures of intergenerational income mobility, including measures of relative and absolute mobility, and more complex measures of “inequality of opportunity”. Our main contribution is to compare rankings across parent-child and grandparent-child correlations. We present our analyses for three statistics, correlations in education and slope coefficients in earnings ranks and log earnings (the IGE).
@@
Moreover, Figure 5 indicates that educational and earnings mobility are not strongly related; compare for example the five largest municipalities that are labeled in the graph in the upper with the lower two panels. Detailed correlations are shown in Table B.2 in the Appendix. Municipalities that have low parent-child educational mobility tend to also have lower mobility in earnings, but this relation is weak (the correlation is 0.06 for fathers and 0.18 for mothers). The relation tends to be even weaker, and in some cases negative, when considering the intergenerational elasticity of earnings (IGE) rather than earnings ranks. This observation stands in contrast to cross-national comparisons, in which educational and income mobility are more strongly related (Blanden 2013, Stuhler 2018). One likely reason is that regional measures of income mobility can be quite noisy (Heidrich, 2017). Indeed, we find that the correlation between educational and income mobility becomes more sizable when restricting our sample to municipalities with larger sample sizes. And interestingly, the relation between education and earnings mobility is generally higher for multi- compared to intergenerational measures (correlations around 0.3 along both the paternal and maternal lines).
@@
Finally, we can compare different measures of earnings mobility. Municipalities characterized by high rank correlations in earnings also have a high IGE. As shown in Table B.2 in the Appendix, this relationship between different earnings measures is similarly strong for inter- and multigenerational correlations. Some of our analyses here are similar to Engzell and Mood (2023), who study the relationships across a multitude of different measures of intergenerational mobility in Sweden using variation over cohorts.
@@
We conclude that inter- and multigenerational mobility measures are closely related, with regions characterized by high parent-child mobility also demonstrating high three-generation mobility. In contrast, education- and earnings-based mobility measures are generally not as close substitutes. Consistent with this conclusion, we observe that the largest Swedish cities exhibit opposing patterns in educational and earnings mobility. While Malmö (the third largest city) is characterized by low earnings mobility, it ranks closer to the Swedish average in terms of educational mobility. Uppsala (the fourth largest city) instead is characterized by very low educational mobility and higher than average earnings mobility. As we saw in Figure 3b, Malmö is an outlier also in terms of absolute upward mobility. The expected earnings rank of children at the 25th percentile of the paternal distribution is 0.43, which places the city in the bottom decile of all municipalities. [^26] An open question that we do not explore further here is whether observed earnings for residents in Malmö might be affected by cross-border commuting to Copenhagen. Still, this rate of upward mobility compares favorably to many regions in the US, in which the expected earnings rank at the 25th percentile can be as low as 0.36 (Chetty et al., 2014).
@@ 5 The Great Gatsby Curve Across Regions
We focus here on the relation between inter- and multigenerational mobility with cross-sectional inequality. A robust pattern in international comparisons is that intergenerational income mobility and inequality are negatively associated across countries (Blanden, 2013; Corak, 2013). In addition, many theories on intergenerational transmission predict that income persistence varies across the income distribution (Durlauf, Kourtellos and Tan, 2022). One particular mechanism contributing to both intergenerational and cross-sectional inequality is the segregation of families into economically homogeneous neighborhoods (Durlauf, 1996).
@@
P25 upward rank mobility $-0.356$ $0.207$ $-0.125$ $-0.048$
@@
P25 upward rank mobility $-0.117$ $0.351$ $0.097$ $0.285$
@@
The first row of Table 3 reports the full-sample equivalent to subfigure a in Figure 6, correlating the rank slope in earnings with the Gini coefficient across all 290 Swedish municipalities. The correlations are close to 0.2 when considering the pooled sample and father-daughter pairs, but lower and insignificant for father-son pairs. [^28] Given the that correlations are constrained within [−1,1] we report p-values rather than standard errors; these p-values do not account for the fact that the underlying statistics are themselves estimated. The correlations are close to zero and insignificant when considering the intergenerational elasticity of earnings (i.e., the log-log regression slope), which might partly stem from that the IGE is mechanically negatively (positively) related to the variance of father’s (child’s) log earnings. [^29] Another approach would be to use the mean of the Gini indices in the father’s and the child’s generations, and thus isolate the analysis from this mechanical relationship (as in e.g. Corak, 2013). One could also standardize log earnings within generations and relate the intergenerational correlation in log earnings with the Gini index. In our data, both these approaches result in positive and sometimes statistically significant relationships (not shown here). The third row shows that, among sons as well as in the pooled sample, the expected earnings rank of those at the 25th percentile of the parental distribution is significantly lower when income inequality is high. Surprisingly, this relation switches sign when only considering daughters, presumably due to female labor force participation being higher in more unequal municipalities. However, apart from this latter result and the results based on the IGE, we tend to confirm the pattern documented in the discourse on the Great Gatsby Curve: intergenerational persistence tends to be systematically higher, and upward mobility significantly lower (for sons), when earnings inequality is high.
@@
This evidence adds to a small number of recent studies that document a similar “Great Gatsby Curve” across regions within countries, although rarely for the three-generation case. [^30] Examples include Chetty et al. (2014) for the U.S. or Güell et al. (2018) for Italy, among many others. For Sweden, Brandén (2019) shows that men who grew up in regions with high levels of income inequality experienced less intergenerational mobility, and highlights the role of educational attainment and cognitive and non-cognitive skills in this relationship. Heidrich (2017) also provides evidence on a “Great Gatsby Curve” across regions within Sweden. We find that the strength of within-country “Great Gatsby Curves” depends on the choice of mobility statistic and how and in what generation inequality is measured. Such choices can also substantially impact the magnitudes of the Gatsby relationship. We further find that the link between intergenerational mobility and income inequality is even more pronounced for educational outcomes than for earnings, which is notable given that education and earnings mobility statistics are only imperfectly related (see Section 4). These patterns are replicated, and the association between income and educational mobility even stronger, in the multigenerational case.

### v2

@@ 3.2 Mobility maps
We can compare our estimates to Heidrich:2017aa, who provides detailed national and regional estimates of intergenerational income mobility for the Swedish population born between 1968 and 1976. We consider more recent cohorts and our empirical specification differs in various aspects. In particular, we predict prime-age earnings based on all available earnings data of each individual, compared to using a conventional shorter-run average of observed earnings or income. Another important difference is that we consider variation on the finer municipality level, while Heidrich focuses on the broader definition of local labor market regions. [^19] One could debate which level of aggregation makes more sense. Child care, compulsory school and high schools are organized at the municipality level, and financed by municipal income taxes. Many other local amenities and services are also best distinguished at the municipal level. Local labor market regions are larger and defined based on commuting patterns. Different divisions are thus likely to capture differences in mechanisms underlying intergenerational transmission to different extents. The exact division is not key in our setting, as we primarily aim at making a set of conceptual and methodological points. Consistent with this and other differences in specification, we find slightly higher earnings persistence. Nevertheless, the spatial pattern appears largely similar. For example, Heidrich:2017aa notes that several local labor markets in the south of Sweden and around Stockholm are doing particularly well in lifting children from lower income families, while absolute upward mobility is particularly low in municipalities in the Swedish inland located close to the Norwegian border. These patterns are also visible in our Appendix Figure A.5b. Given that we study different birth cohorts, this similarity also implies that mobility differences between regions are systematic and at least somewhat persistent, rather than just temporary.
@@ 3.3 Linear approximations
Regional comparisons of intergenerational mobility, such as those reported in Figures 1 and 2, are subject to two methodological concerns. The first is whether intergenerational relations can be sufficiently well summarized by simple summary measures, such as linear regression or correlation coefficients. CholliDurlauf2022 note that theoretical models usually imply nonlinear relationships between parent and child status, which are often ignored in practice.
For illustration, Figure 3a plots the mean child earnings rank against the parental earnings rank at the national level. The rank-rank relation is linear over most of the distribution, but much steeper at the very bottom and top. [^20] NybomStuhlerJHR2017 document a similarly steep relation at the very bottom of the parental income distribution. The conditional expectation function (CEF) is therefore not fully captured by a linear regression, in contrast to the pattern for the U.S. as described by CHKS_Where_QJE. However, it serves as a good approximation over most of the distribution, such as between the 10th and 90th percentiles. The expectation of child years of schooling conditional on father’s years of schooling is approximately linear for most of the distribution as well (see Appendix Figure A.6).
@@
While the CEF appears sufficiently linear on the national level, it does not necessarily follow that the conditional expectations are equally linear on the regional level. For illustration, Figure 3b plots the mean child earnings rank conditional on parent rank deciles for Sweden’s five largest municipalities. Upward mobility is substantially lower, and the rank-rank slope steeper, in Malmö (diamonds) compared to for example Uppsala (grey circles). However, apart from such differences for the bottom decile, the relationships are approximately linear in all cities; while it is still somewhat steeper at the bottom and the top of the parental earnings distribution, this pattern is less pronounced at the regional level. We provide the corresponding graphs for the grandfather-child relationship in Figure A.7 in the Appendix. The region-specific multigenerational CEFs are approximately linear, and less steep at the bottom of the income distribution than their national-level counterpart. [^21] Further, we conducted a simple analysis for all municipalities by comparing the regression $R^{2}$ from municipality-specific linear and quadratic regressions. The quadratic model generates an $R^{2}$ that on average is only 2.7% larger than the mean $R^{2}$ from the linear model. This number drops to only 1.4% if we weight the averages by number of individuals rather than by municipality. Linear measures therefore appear well suited for cross-regional comparisons.
@@ 3.4 Sampling error
A second concern in fine-grained regional comparisons is sample size: Some of the variation across municipalities might not be “real” variation but reflect sampling error, in particular for smaller municipalities for which comparatively few parent-child pairs are observed. [^22] As we observe (almost) the entire population, our estimates do not suffer from traditional sampling error. However, the observed mobility for a given set of parent-child pairs can be interpreted as a random draw from the “potential” distribution for that municipality. While the mean number of father-child pairs per municipality for children born between 1981 and 1989 is 3,183 for education and 3,099 for earnings, in 57 of the municipalities (59 for earnings) we observe fewer than 1,000 pairs. [^23] Moreover, inference based on earnings ranks is complicated by the fact that each unit’s rank depends on which other units have been sampled (MogstadWilhelm23). We also observe fewer grandfather-child pairs than father-child pairs, and especially so for earnings. Education is observed for all who have resided in Sweden at some point as adults, while earnings is only observed for those who at some point from 1968 have recorded taxable labor earnings above a certain lower amount (see Section 2). To probe the influence of sampling error, we implement two tests.
@@
Notes: Density plot of the estimated intergenerational correlations in years of schooling for father-child pairs across 160 “small” municipalities ($\leq$ 2,000 observations), 130 “large” municipalities (>> 2,000 observations), and corresponding “placebo” municipalities in which we randomly allocated pairs across municipalities.
@@
First, we split the sample into smaller ($\leq$ 2,000 observations) and larger municipalities (>> 2,000 observations), with the expectation that the influence of sampling error should be more limited for the latter. Figure 4 plots the resulting densities of the father-child correlation in years of schooling. The correlations vary nearly as much for large as for small municipalities (the standard deviations are 0.038 vs. 0.048). Figure A.8 in the Appendix shows corresponding distributions for earnings. [^24] Figure 4 shows that the educational correlations are on average substantially lower for small municipalities, a pattern that is possibly related to the greater homogeneity in educational attainment in smaller municipalities. And while the estimated coefficient in an OLS regression of the municipal parent-child correlation on sample size is large and significant, it loses significance if we control for the standard deviation of years of schooling. The difference in means between small and large municipalities for earnings ranks is smaller, but qualitatively similar (see Figure A.8 in the Appendix). Second, we compute “placebo correlations” by reshuffling parent-child pairs randomly across all 290 municipalities, such that variation in the placebo correlations solely reflects chance. Figure 4 shows that for both small and large municipalities, the resulting density is much narrower than the actual variation of intergenerational correlations across municipalities. Moreover, there is considerably less dispersion in the placebo distribution for large than for small municipalities – the standard deviation of the former is less than half of the one for the latter. [^25] Considering instead changes, the decrease in dispersion is larger for large than for small municipalities when comparing the placebo with the true distributions, with a drop in the std. dev. of about 70% compared to 45% for small municipalities. These differences between small and large municipalities are even more pronounced for earnings ranks. In particular, the std. dev. of the placebo distribution of earnings rank slopes of large municipalities is about 44% of the one for small ones.
@@
Still, some of the reported variation between municipalities reflects noise, in particular for small municipalities, and sampling error might explain much of the regional variation in studies that use less data than we do here. Sampling error may have a greater effect for earnings-based measures, as earnings are themselves subject to transitory noise (mazumder2005fortunate). This is the case in our setting as well, even though we use comparatively high-quality earnings measures (see Appendix Figure A.8). Few studies have addressed this issue systematically. Rather than estimating separate regressions for each region, Heidrich:2017aa estimates a multilevel model that explicitly accounts for the influence of sampling error on the variation of regional mobility estimates. This approach effectively “shrinks” the region-specific estimates to their mean, in particular for small regions with imprecise estimates. Risa:2019aa implements a similar shrinkage procedure for mobility estimates for Norway. In addition, Heidrich:2017aa reports only those estimates that are significantly differently from the mean, replacing all other region-specific estimates with that mean.
@@ 4 Are Regional Rankings Stable?
Our main focus is on whether intergenerational (parent-child) associations – which can be estimated in many settings – reflect similar underlying mobility processes as more data-demanding multigenerational associations. This analysis is complementary to a recent analysis by deutscher2023measuring, who compare the ranking of Australian regions across different measures of intergenerational income mobility, including measures of relative and absolute mobility, and more complex measures of “inequality of opportunity”. Our main contribution is to compare rankings across parent-child and grandparent-child correlations. We present our analyses for three statistics, correlations in education and slope coefficients in earnings ranks and log earnings (the IGE).
@@
Moreover, Figure 5 indicates that educational and earnings mobility are not strongly related; compare for example the five largest municipalities that are labeled in the graph in the upper with the lower two panels. Detailed correlations are shown in Table B.1 in the Appendix. Municipalities that have low parent-child educational mobility tend to also have lower mobility in earnings, but this relation is weak (the correlation is 0.06 for fathers and 0.18 for mothers). The relation tends to be even weaker, and in some cases negative, when considering the intergenerational elasticity of earnings (IGE) rather than earnings ranks. This observation stands in contrast to cross-national comparisons, in which educational and income mobility are more strongly related (Blanden2011Survey, Stuhler2018JRC). One likely reason is that regional measures of income mobility can be quite noisy (Heidrich:2017aa). Indeed, we find that the correlation between educational and income mobility becomes more sizable when restricting our sample to municipalities with larger sample sizes. And interestingly, the relation between education and earnings mobility is generally higher for multi- compared to intergenerational measures (correlations around 0.3 along both the paternal and maternal lines).
@@
Finally, we can compare different measures of earnings mobility. Municipalities characterized by high rank correlations in earnings also have a high IGE. As shown in Table B.1 in the Appendix, this relationship between different earnings measures is similarly strong for inter- and multigenerational correlations. Some of our analyses here are similar to engzell2023understanding, who study the relationships across a multitude of different measures of intergenerational mobility in Sweden using variation over cohorts.
@@
We conclude that inter- and multigenerational mobility measures are closely related, with regions characterized by high parent-child mobility also demonstrating high three-generation mobility. In contrast, education- and earnings-based mobility measures are generally not as close substitutes. Consistent with this conclusion, we observe that the largest Swedish cities exhibit opposing patterns in educational and earnings mobility. While Malmö (the third largest city) is characterized by low earnings mobility, it ranks closer to the Swedish average in terms of educational mobility. Uppsala (the fourth largest city) instead is characterized by very low educational mobility and higher than average earnings mobility. As we saw in Figure 3b, Malmö is an outlier also in terms of absolute upward mobility. The expected earnings rank of children at the 25th percentile of the paternal distribution is 0.43, which places the city in the bottom decile of all municipalities. [^26] An open question that we do not explore further here is whether observed earnings for residents in Malmö might be affected by cross-border commuting to Copenhagen. Still, this rate of upward mobility compares favorably to many regions in the US, in which the expected earnings rank at the 25th percentile can be as low as 0.36 (CHKS_Where_QJE).
@@ 5 The Great Gatsby Curve Across Regions
We focus here on the relation between inter- and multigenerational mobility with cross-sectional inequality. A robust pattern in international comparisons is that intergenerational income mobility and inequality are negatively associated across countries (Blanden2011Survey; CorakJEP2013). In addition, many theories on intergenerational transmission predict that income persistence varies across the income distribution (durlauf2022great). One particular mechanism contributing to both intergenerational and cross-sectional inequality is the segregation of families into economically homogeneous neighborhoods (Durlauf1996).
@@
P25 upward rank mobility 0.356-0.356 $0.207$ 0.125-0.125 0.048-0.048
@@
P25 upward rank mobility 0.117-0.117 $0.351$ $0.097$ $0.285$
@@
The first row of Table 3 reports the full-sample equivalent to subfigure a in Figure 6, correlating the rank slope in earnings with the Gini coefficient across all 290 Swedish municipalities. The correlations are close to 0.2 when considering the pooled sample and father-daughter pairs, but lower and insignificant for father-son pairs. [^28] Given the that correlations are constrained within [−1,1] we report p-values rather than standard errors; these p-values do not account for the fact that the underlying statistics are themselves estimated. The correlations are close to zero and insignificant when considering the intergenerational elasticity of earnings (i.e., the log-log regression slope), which might partly stem from that the IGE is mechanically negatively (positively) related to the variance of father’s (child’s) log earnings. [^29] Another approach would be to use the mean of the Gini indices in the father’s and the child’s generations, and thus isolate the analysis from this mechanical relationship (as in e.g. CorakJEP2013). One could also standardize log earnings within generations and relate the intergenerational correlation in log earnings with the Gini index. In our data, both these approaches result in positive and sometimes statistically significant relationships (not shown here). The third row shows that, among sons as well as in the pooled sample, the expected earnings rank of those at the 25th percentile of the parental distribution is significantly lower when income inequality is high. Surprisingly, this relation switches sign when only considering daughters, presumably due to female labor force participation being higher in more unequal municipalities. However, apart from this latter result and the results based on the IGE, we tend to confirm the pattern documented in the discourse on the Great Gatsby Curve: intergenerational persistence tends to be systematically higher, and upward mobility significantly lower (for sons), when earnings inequality is high.
@@
This evidence adds to a small number of recent studies that document a similar “Great Gatsby Curve” across regions within countries, although rarely for the three-generation case. [^30] Examples include CHKS_Where_QJE for the U.S. or GuellPellizzariPicaMora2018 for Italy, among many others. For Sweden, Branden2019 shows that men who grew up in regions with high levels of income inequality experienced less intergenerational mobility, and highlights the role of educational attainment and cognitive and non-cognitive skills in this relationship. Heidrich:2017aa also provides evidence on a “Great Gatsby Curve” across regions within Sweden. We find that the strength of within-country “Great Gatsby Curves” depends on the choice of mobility statistic and how and in what generation inequality is measured. Such choices can also substantially impact the magnitudes of the Gatsby relationship. We further find that the link between intergenerational mobility and income inequality is even more pronounced for educational outcomes than for earnings, which is notable given that education and earnings mobility statistics are only imperfectly related (see Section 4). These patterns are replicated, and the association between income and educational mobility even stronger, in the multigenerational case.